Let X be a binomial random variable with p = 0.1 and n = 10. Calculate the following probabilities.

(a) P(X ≤ 2) 

(b) P(X > 8)

(c) P(X = 4) 

(d) P(5 ≤ X ≤ 7)


1
Expert's answer
2021-06-28T02:15:23-0400

(a)

"n=10, p=0.1, q=1-p=1-0.1=0.9"


"P(X\leq 2)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{10}{0}(0.1)^0(0.9)^{10-0}+\dbinom{10}{1}(0.1)^1(0.9)^{10-1}"

"+\dbinom{10}{2}(0.1)^2(0.9)^{10-2}=0.3486784401"

"+0.387420489+0.1937102445"

"=0.9298091736"

(b)

"n=10, p=0.1, q=1-p=1-0.1=0.9"

"P(X>8)=P(X=9)+P(X=10)"

"=\dbinom{10}{9}(0.1)^9(0.9)^{10-9}+\dbinom{10}{10}(0.1)^{10}(0.9)^{10-10}"

"=0.000000009+0.0000000001"

"=0.0000000091"



(c)

"n=10, p=0.1, q=1-p=1-0.1=0.9"

"P(X=4)=\dbinom{10}{4}(0.1)^4(0.9)^{10-4}=0.011160261"

(d)

"n=10, p=0.1, q=1-p=1-0.1=0.9"


"P(5\leq X\leq 7)=P(X=5)+P(X=6)+P(X=7)"

"=\dbinom{10}{5}(0.1)^5(0.9)^{10-5}+\dbinom{10}{6}(0.1)^6(0.9)^{10-6}"

"+\dbinom{10}{7}(0.1)^7(0.9)^{10-7}=0.0014880348"

"+0.000137781+0.000008748"

"=0.0016345638"